# Agentic AI for Software: thoughts from Software Engineering community

### v2

@@ Abstract
Conceptually, the key to successfully developing trustworthy agentic AI-based software workflows will be to resolve the core difficulty in software engineering - the deciphering and clarification of developer intent. Specification inference, or deciphering the intent, thus lies at the heart of many software tasks, including software maintenance and program repair. A successful deployment of agentic technology into software engineering would involve making conceptual progress in such intent inference via agents. We discuss, to some length, the AutoCodeRover agent which embodies such intent inference. The agent has been integrated into the SonarQube static analysis tool, which is used by thousands of enterprise customers.
@@
Since the 1970’s when programming in modern existing programming languages firmly took root - the focus has been on programming in the large. The start of this period can be traced back to around fifty years ago - when the first Hello World program was written by Brian Kerninghan around 1972-3 in the B programming language, which is the predecessor of C. Since the writing of the first ”Hello World” program and the initial seminal innovations there have been only two really major shifts in the software industry affecting both the technology as well as the business of software.
@@
In 2025, we are now at the second major shift which affects both the software technology and business- the arrival of agentic AI. With the use of agentic AI in software - the change is beyond the hosting of the software - instead of the engineering of the software itself is automated to a large degree. Since there is greater automation - this eases the writing of code fragments capturing different functionality. However the correctness of the code and the overall trust in the agent output becomes a key concern that needs addressing. Thus, perhaps for the first time since the writing of the ”Hello World” program - we face the central issue of programming with trust, instead of programming in the large.
@@
Figure 1: Arrival of Agentic AI in Software - five decades after ”Hello World”.
@@
AutoCodeRover (Zhang et al. 2024) gives a different agentic AI approach by involving program analysis. Given a software issue, its core intuition is to extract specifications (of intended behavior) from the program structure. This is achieved via code search. Starting from the code elements (classes, methods, APIs) mentioned in the issue - it progressively searches other code elements to perform fault localization - finding where the code changes are needed. This is then followed by a patch generation stage which considers the knowledge/specifications gained from fault localization. A follow-up work SpecRover (Ruan, Zhang, and Roychoudhury 2025) makes the implicit intent inference in AutoCodeRover more explicit by inferring the intended behavior of each of the functions (or units off code) identified by fault localization. Figure 3 shows a simple example of issue resolution via intent inference. The issue being resolved is shown at the top. The code-search starts from the ModelChoiceField class which is mentioned in the natural language issue. It then progresses by looking up the signature of this class - and then further looking into the implementation of the methdods. The AutoCodeRover agent thus iteratively refines its lower-level understanding of the intended program behavior (which is always described at a high level by the issue), from the program structure such as classes / methods.
@@
What is the benefit of the agent working on program representations? The main benefit is the inference of intended program behavior by the agent which can be used to provide explanations of the agent’s actions. It enhances trust in the coding agent, and allows one to employ agents in more critical scenarios (that demand greater trust) than vibe-coding. We can also observe this from the current real-life usage of AutoCodeRover, which has now been integrated into the widely used SonarQube tool (Gaudin and Mallet 2010). SonarQube is used by thousands of enterprises for code quality and security - it provides inspection of bugs, code smells and vulnerabilities. Such bug reports generated by SonarQube can be treated as issues which are then automatically fixed by AutoCodeRover. This use-case shows the promise of agentic AI in providing production grade tool support for critical code.
@@
Issue resolution forms the core of program improvement and maintenance of a large code-base. While issue resolution and prompt based code generation have captured significant attention - the opportunities of using agentic AI for code are by no means restricted to these areas. With the growth of automation in software development, there will arguably be an explosion of AI-generated code. This will exacerbate the issue of verification and validation (V & V) of the generated code. Furthermore, as the generated code gets integrated to code-bases, the interplay of legacy human written code and the generated code - will lead to subtle errors and vulnerabilities. Even if the agentic AI that produces code (fragments) is trustworthy - the integration of AI generated code can create new trust issues in software projects further down the line. There is a long history of prominent vulnerabilities caused by (human written) software. The latest prominent example is perhaps CrowdStrike outage in 2024, where the security system itself had a bug that caused worldwide outages. At the same time, AI generated code from prompts have been reported to have many vulnerabilities as per various academic studied (see e.g., (Pearce et al. 2025)). There may be a lingering ”trust deficit” in AI generated code.
Looking forward, one prominent direction thus could be to put agentic AI itself to use in combating the trust deficit from AI generated code ! Thus, given a key critical high level natural language requirement - agentic AI may help clarify the requirement by understanding its intent, as well as understanding how it is implemented at a low level in the code-base. This may lead to capabilities of security audits which have been recently studied (Guo et al. 2025). In fact, it can go beyond audits by flagging violations of the policy and fixing these violations via code changes albeit by an issue resolution in the back-end.
@@
Last but not the least, agentic AI can be put to use for formal verification of code (Cadar and Roychoudhury 2025), which will be possible if the AI agents can interpret representations of formal proofs generated by theorem provers. Just like agents like AutoCodeRover can interpret program representations and analyze them, we can devise agents which can interpret formal proof representations, analyze them, and communicate with a theorem prover via an iterative refinement loop to ”improve” the proof. This will also contrast with approaches like Alphaproof which are for proving general mathematical theorems. Instead, a program verification agent can again leverage the program structure like classes / methods to structure the proof and its intermediate lemmas. This remains an attractive future direction.

### v3

@@ Abstract
Conceptually, the key to successfully developing trustworthy agentic AI-based software workflows will be to resolve the core difficulty in software engineering - the deciphering and clarification of developer intent. Specification inference, or deciphering the intent, thus lies at the heart of many software tasks, including software maintenance and program repair. A successful deployment of agentic technology into software engineering would involve making conceptual progress in such intent inference via agents. We discuss, to some length, the AutoCodeRover agent which embodies such intent inference. The agent has been integrated into the SonarQube static analysis tool.
@@
Since the 1970’s when programming in modern existing programming languages firmly took root - the focus has been on programming in the large. The start of this period can be traced back to around fifty years ago - when the first Hello World program was written by Brian Kerninghan in 1972-3 in the B programming language, which is the predecessor of C. Since the writing of the first ”Hello World” program and the initial seminal innovations, there have been only two really major shifts in the software industry affecting both the technology as well as the business of software.
@@
In 2025, we are now at the second major shift which affects both the software technology and business- the arrival of agentic AI. With the use of agentic AI in software - the change is beyond the hosting of the software - instead, the engineering of the software itself is automated to a large degree. Since there is greater automation - this eases the writing of code fragments capturing different functionality. However, the correctness of the code and the overall trust in the agent output becomes a key concern that needs addressing. Thus, perhaps for the first time since the writing of the ”Hello World” program - we face the central issue of programming with trust, instead of programming in the large.
@@
Figure 1: Arrival of Agentic AI in Software - more than fifty years after ”Hello World”.
@@
AutoCodeRover (Zhang et al. 2024) gives a different agentic AI approach by involving program analysis. Given a software issue, its core intuition is to extract specifications (of intended behavior) from the program structure. This is achieved via code search. Starting from the code elements (classes, methods, APIs) mentioned in the issue - it progressively searches other code elements to perform fault localization - finding where the code changes are needed. This is then followed by a patch generation stage which considers the knowledge/specifications gained from fault localization. A follow-up work SpecRover (Ruan, Zhang, and Roychoudhury 2025) makes the implicit intent inference in AutoCodeRover more explicit by inferring the intended behavior of each of the functions (or units off code) identified by fault localization. Figure 3 shows a simple example of issue resolution via intent inference. The issue being resolved is shown at the top. The code-search starts from the ModelChoiceField class which is mentioned in the natural language issue. It then progresses by looking up the signature of this class - and then further looking into the implementation of the methdods. The AutoCodeRover agent thus iteratively refines its lower-level understanding of the intended program behavior from the program structure such as classes / methods.
@@
What is the benefit of the agent working on program representations? The main benefit is the inference of intended program behavior by the agent which can be used to provide explanations of the agent’s actions. It enhances trust in the coding agent, and allows one to employ agents in more critical scenarios (that demand greater trust) than vibe-coding. We can also observe this from the current real-life usage of AutoCodeRover, which has now been integrated into the widely used SonarQube tool (Gaudin and Mallet 2010). SonarQube is used by enterprises for code quality and security - it provides inspection of bugs, code smells and vulnerabilities. Such bug reports generated by SonarQube can be treated as issues which are then automatically fixed by AutoCodeRover. This shows the promise of agentic AI in providing support for critical code.
@@
Issue resolution forms the core of program improvement and maintenance of a software project. While issue resolution and prompt based code generation have captured significant attention - the opportunities of using agentic AI for code are by no means restricted to these areas. With the growth of automation in software development, there will arguably be an explosion of AI-generated code. This will exacerbate the issue of verification and validation (V & V) of the generated code. Furthermore, as the generated code gets integrated to code-bases, the interplay of legacy human written code and the generated code - will lead to subtle errors and vulnerabilities. Even if the agentic AI that produces code (fragments) is trustworthy - the integration of AI generated code can create new trust issues in software projects further down the line. There is a long history of prominent vulnerabilities caused by (human written) software. The latest prominent example is perhaps CrowdStrike outage in 2024, where the security system itself had a bug that caused worldwide outages. At the same time, AI generated code from prompts have been reported to have many vulnerabilities (see e.g., (Pearce et al. 2025)). There may be a lingering ”trust deficit” in AI generated code.
Looking forward, one prominent direction thus could be to put agentic AI itself to use in combating the trust deficit from AI generated code ! Thus, given a key critical high level natural language requirement - agentic AI may help clarify the requirement by understanding its intent, as well as understanding how it is implemented at a low level in the code-base. This may lead to capabilities of security audits which have been recently studied (Guo et al. 2025). In fact, it can go beyond audits by flagging violations of the policy and fixing these violations via code changes albeit by an issue resolution agent in the back-end.
@@
Last but not the least, agentic AI can be put to use for formal verification of code (Cadar and Roychoudhury 2025). Just like agents like AutoCodeRover can interpret program representations and analyze them, we can devise AI agents which can interpret formal proof representations, analyze them, and communicate with a theorem prover via an iterative refinement loop to ”improve” the proof. This will also contrast with approaches like Alphaproof [^1] https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/ which help prove general mathematical theorems. Instead, a program verification agent can again leverage the program structure like classes / methods to structure the proof and its intermediate lemmas. This remains an attractive future direction.